Roland's roller derby car traveled 4 1/2 feet past the bottom of the ramp. Elenor's car traveled 6 3/4 feet past the ramp. How much farther did Eleanor's car travel?

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

Roland's roller derby car traveled 4 1/2 feet past the bottom of the ramp. Elenor's car traveled 6 3/4 feet past the ramp. How much farther did Eleanor's car travel?
------
Ans: 6 3/4 - 4 1/2 = 27/4 - 9/2 = 27/4 - 18/4 = 9/4 = 2 1/4 feet
=================================
Cheers,
Stan H.

Answer by Alan3354(69443)

That's soapbox derby, not roller derby.